OpenStudy (anonymous):

What is the sum of a 16 term sequence where the last term is 21and the difference is -6?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@lgbasallote ? This is different from my other question so I'm not understanding

OpenStudy (lgbasallote):

yeah...this takes more.. first you need to find the first term

OpenStudy (lgbasallote):

\[a_n = a_1 + (n-1)d\] remember i said a_n is the last term? so in this case it would be 21 then common difference is -6 number of terms is 16 \[21 = a_1 + (16-1)(-6)\] find a_1

OpenStudy (anonymous):

a1 is the 1st term so how am I supposed to find it? Am I just suppose to simplify that? @lgbasallote

OpenStudy (lgbasallote):

yes. and then isolate a_1 just like in usual algebra

OpenStudy (anonymous):

I got 21= a1+ (-90) What next?

OpenStudy (lgbasallote):

next you add 90 to both sides

OpenStudy (anonymous):

a1= 111?

OpenStudy (lgbasallote):

right. so now that's the first term. now use \[S_n = \frac{n(a_n + a_1)}{2}\] to get the sum

OpenStudy (anonymous):

What's Sn?

OpenStudy (lgbasallote):

sum of n terms

OpenStudy (lgbasallote):

that's what you're looking for

OpenStudy (lgbasallote):

just substitute n, a1 and an.. dont worry about the notations

OpenStudy (anonymous):

I don't understand how to add a_21+111.

OpenStudy (lgbasallote):

a_21? where did that come from?

OpenStudy (anonymous):

the number of terms? I'm very confused can you solve the end so i can see what you did.

OpenStudy (lgbasallote):

\[\large \implies \frac{16(21 + 111)}{2}\] remember what i said above n = 16 <--because there are 16 terms a_n =21 <--because it's the last term a_1 = 111<--you solved for this

OpenStudy (anonymous):

Ohhhhh ok!

OpenStudy (anonymous):

I got 1056

OpenStudy (anonymous):

So 1,056 and 111 are both the answers?

OpenStudy (lgbasallote):

111 is not an answer...it's just looking for the sum..not the first term...so 1056 is the only answer`

OpenStudy (anonymous):

Thank youu

OpenStudy (lgbasallote):

welcome